# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.3 Part $M\subset M_{0}$
$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
@@
×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{(\tau-1)/2}),$ (A.33)
@@
$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})+O_{p}(T^{\tau})+O_{p}(T^{(\tau-1)/2})$
$=O_{p}(T^{\tau}),$ (A.34)
which is the dominant order. This implies that
$T^{-\tau}(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}})$ $=T^{-\tau}\mathbf{III}_{d}+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-(\tau+1)/2}),$ (A.35)
@@
which converges to a positive semi-definite matrix. However, note how by using the fact that $\overline{\mathbf{Q}}_{M_{0}}=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})$, we have that
$T^{-\tau}\overline{\mathbf{Q}}_{M_{0}}$ $=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=T^{-\tau}\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{X}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{T^{1+\tau}}\frac{1}{N}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}-T^{-1}\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}+O_{p}(\xi_{N,T,\tau}^{-1})$
$=O_{p}(T^{-\tau})+O_{p}(\xi_{N,T,\tau}^{-1})+O_{p}(T^{-1})$
$=o_{p}(1),$ (A.36)
because the idiosyncratics are stationary. Combining (6.3) and (6.3), we get
$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+o(1)$
↛$\not\to_{p}c>0,$ (A.37)
as in total $T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right](T^{-\tau}\overline{\mathbf{Q}}_{M_{0}})^{-1}$ does not converge to a positive semi-definite matrix. Therefore
$\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)\not\to 0,$ (A.38)
and the under-specification risk is not eliminated.
6.4 Part $M_{0}\subset M$

### v3

@@ 5.3.1 Case of $M\subset M_{0}$
$\leq T^{-(\tau+1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
@@
$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{-(\tau+1)/2})=o_{p}(T^{-\tau}),$ (A.31)
@@
$T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$ $=T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-T^{-\tau}\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$+T^{-\tau}\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$=\mathbf{III}_{d}+o_{p}(T^{-\tau}),$ (A.32)
@@
which is a positive semi-definite matrix and $o_{p}(T^{-\tau})=O_{p}(\xi_{N,T,\tau}^{-1})$. Further, note how by using the fact that $\overline{\mathbf{Q}}_{M_{0}}=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+o_{p}(1)$ from (A.23), we have that
$\overline{\mathbf{Q}}_{M_{0}}$ $=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}+o_{p}(1)$
$=\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}}+o_{p}(1)$
$=\frac{1}{T}\frac{1}{N}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{X}_{i}+o_{p}(1)$
$=\frac{1}{T}\frac{1}{N}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}}\mathbf{V}_{i}+o_{p}(1)$
$=\frac{1}{N}\sum_{i=1}^{N}T^{-1}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}-T^{-1}\frac{1}{N}\sum_{i=1}^{N}T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}})^{+}T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}+o_{p}(1)$
$=\frac{1}{N}\sum_{i=1}^{N}T^{-1}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}+o_{p}(1)$
$\to_{p}\boldsymbol{\Sigma}_{\mathbf{v}},$ (A.33)
because the idiosyncratics are stationary. This is a positive definite matrix. In what follows, we use Theorem 2 and Exercise 1 in Section 4 of Magnus and Neudecker, (2019). In particular,
$\ln\mathrm{det}(\mathbf{A}+\mathbf{H})=\ln\mathrm{det}(\mathbf{A})+\mathrm{tr}(\mathbf{A}^{-1}\mathbf{H})+O\left(\left\|\mathbf{A}^{-1}\mathbf{H}\right\|^{2}\right),$ (A.34)
for a positive definite matrix $\mathbf{A}$ and a symmetric perturbation $\mathbf{H}$. Further, to simplify notation, let $\boldsymbol{\Omega}_{M,M_{0}}:=T^{-\tau}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$ and $\boldsymbol{\Omega}_{M,M_{0}}^{0}:=\mathbf{III}_{d}\times\overline{\mathbf{Q}}_{M_{0}}^{-1}$, which we know is a positive semi-definite matrix. Next, let $\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}=\boldsymbol{\Omega}_{M,M_{0}}-\boldsymbol{\Omega}_{M,M_{0}}^{0}$, where we know that $\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau})$. In our setting, $\mathbf{A}=\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}$ and $\mathbf{H}=\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}$. Therefore,
$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)$
$+\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)$
$-\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]$
$+O_{p}\left(\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|^{2}\right)$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}).$ (A.35)
Here, we used the following important facts. Let $\lambda_{NT,j}\geq 0$ be an eigenvalue of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ for $j=1,\ldots,k$ with the probability limit $\lambda>0$ or 0 immediately. As this positive semi-definite matrix is nonzero even in the limit, we know that not all eigenvalues are zero. Then we know that the eigenvalues of $\mathbf{B}=\mathbf{I}_{k}+c\mathbf{A}$ are given by $1+c\lambda_{j}$ for a constant $c$, where $\lambda_{j}$ is an eigenvalue of generic square matrix $\mathbf{A}$. In particular, for the $j$-th eigenvector of $\mathbf{A}$ ($\boldsymbol{\nu}_{j}$), we have
$\mathbf{B}\boldsymbol{\nu}_{j}=(\mathbf{I}_{k}+c\mathbf{A})\boldsymbol{\nu}_{j}=\boldsymbol{\nu}_{j}+c\mathbf{A}\boldsymbol{\nu}_{j}=1\times\boldsymbol{\nu}+c\lambda_{j}\boldsymbol{\nu}_{j}=(1+c\boldsymbol{\lambda}_{j})\boldsymbol{\nu}_{j}.$ (A.36)
Next, we know that eigenvalues of $\mathbf{B}^{-1}=(\mathbf{I}_{k}+c\mathbf{A})^{-1}$ are given by $\frac{1}{1+c\lambda_{j}}$. Which means that eigenvalues of $\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}$ are given by $\frac{1}{1+T^{\tau}\lambda_{NT,j}}$. Note that we can diagonalize this matrix and obtain
$\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}=\boldsymbol{\Upsilon}\left(\underbrace{\mathrm{diag}\left[\frac{1}{1+T^{\tau}\lambda_{NT,1}},\ldots,\frac{1}{1+T^{\tau}\lambda_{NT,k}}\right]}_{=\boldsymbol{\Lambda}_{N,T}}\right)\boldsymbol{\Upsilon}^{\prime},$ (A.37)
where $\boldsymbol{\Upsilon}$ is an orthonormal vector. Let $b$ denote the number of eigenvalues of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ that is strictly positive. Then kbk-b is zero eigenvalues, which means that we will have kbk-b eigenvalues of $\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}$ that are equal to 1. Then by using the trace properties, we obtain
$\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}\boldsymbol{\Upsilon}^{\prime}\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}\boldsymbol{\Upsilon}^{\prime}\right)}$ $=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}^{2}\boldsymbol{\Upsilon}^{\prime}\right)}$
$=\sqrt{\mathrm{tr}\left(\boldsymbol{\Upsilon}^{\prime}\boldsymbol{\Upsilon}\boldsymbol{\Lambda}_{N,T}^{2}\right)}$
$=\sqrt{\sum_{j=1}^{k}\frac{1}{(1+T^{\tau}\lambda_{NT,j})^{2}}}$
$\to_{p}\sqrt{k-b},$ (A.38)
because $\lambda_{NT,j}=0$ or $\lambda_{NT,j}\to_{p}\lambda_{j}>0$, where the latter case produces $b$ zero entries into the total sum, and the former case gives kbk-b units. Therefore, $\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|=O_{p}(1)$. Further,
$\left\|\mathrm{tr}\left[\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right]\right\|$ $\leq\left\|\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)^{-1}\right\|T^{\tau}\left\|\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right\|$
$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$
$=R_{N,T}.$ (A.39)
Therefore, combining (5.3.1) and (5.3.1), we get
$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\left(\boldsymbol{\Omega}_{M,M_{0}}^{0}+\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}+T^{\tau}\mathrm{d}\boldsymbol{\Omega}_{M,M_{0}}\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)+k(g-m)p_{N,T}+R_{N,T}.$ (A.40)
Clearly, the difference diverges to $\infty$, because $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ converges to a nonzero positive semi-definite matrix. To complete the proof, it is useful to derive the rate at which the divergence occurs. Again, recall that $\lambda_{NT,j}\geq 0$ is an eigenvalue of $\boldsymbol{\Omega}_{M,M_{0}}^{0}$ for $j=1,\ldots,k$ with the probability limit $\lambda>0$ or zero immediately. As this positive semi-definite matrix is nonzero even in the limit, we know that not all eigenvalues are zero. Recall that $b$ is the number of positive eigenvalues. Therefore, using $\ln(1+x)=x+O(x^{2})$ for a small $x$ and $p_{N,T}/\ln(T)=o(1)$, we obtain the following:
$(\mathrm{IC}^{DVS}(M)$ $-\mathrm{IC}^{DVS}(M_{0}))/\ln(T)=\left[\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\boldsymbol{\Omega}_{M,M_{0}}^{0}\right)\right]/\ln(T)+o_{p}(1)$
$=\left[\ln\prod_{j:\lambda>0}(1+T^{\tau}\lambda_{NT,j})\right]/\ln(T)+o_{p}(1)$
$=\left[\sum_{j:\lambda>0}\ln(1+T^{\tau}\lambda_{NT,j})\right]/\ln(T)+o_{p}(1)$
$=\left[\sum_{j:\lambda>0}\ln(T^{\tau}\lambda_{NT,j})\right]/\ln(T)+\left[\sum_{j:\lambda>0}[\ln(1+T^{\tau}\lambda_{NT,j})-\ln(T^{\tau}\lambda_{NT,j})]\right]/\ln(T)+o_{p}(1)$
$=\frac{\tau b\ln(T)}{\ln(T)}+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln(\lambda_{NT,j})+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln\left(1+\frac{1}{T^{\tau}\lambda_{NT,j}}\right)+o_{p}(1)$
$=\tau b+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\ln(\lambda_{NT,j})+\frac{1}{\ln(T)}\sum_{j:\lambda>0}\left(\frac{1}{T^{\tau}\lambda_{NT,j}}+O(T^{-2\tau})\right)+o_{p}(1)$
$\to_{p}\tau b>0,$ (A.41)
because $\lambda_{NT,j}\to_{p}\lambda_{j}>0$ as $(N,T)\to\infty$. This overall implies that
$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(\frac{\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.42)
as $(N,T)\to\infty$.
5.3.2 Case of $M_{0}\subset M$